OpenStudy (anonymous):

n(A)=10, n(AB)=3, and n(A U B)=19, what is n(B)?

OpenStudy (anonymous):

N(A or B) = N(A) + N(B) - N(A and B)

OpenStudy (anonymous):

So you have 19=10+N(B)-3, and you get N(B)=12.